OpenStudy (anonymous):

whats the surface area of a cone with a slant height of 13 and a radius of 13 a.1061.3 b.1075.4 c.1062.9 d.1060.0

OpenStudy (compassionate):

SA = (pi)rs + (pi)^2 1281.8

OpenStudy (anonymous):

nope. these are the options to my question.. 1061.3 1075.4 1062.9 1060.0

OpenStudy (compassionate):

r = 13 m h = 0 m s = 13 m V = 0 m3 L = 530.929 m2 B = 530.929 m2 A = 1061.86 m2

OpenStudy (anonymous):

\(13^2\pi+13^2\pi=338\pi\)

OpenStudy (anonymous):

Of your answers, the one they're likely looking for is 1061.3, which is obtained if you use 3.14 for pi.

OpenStudy (anonymous):

thank you nbouscal & compassionate (: